OpenStudy (anonymous):

prove that-: 2+2=5 ?

OpenStudy (anonymous):

this is wrong question

terenzreignz (terenzreignz):

@03453660 is right, but let's have some fun, anyway. Suppose a = b, where ab is not equal to 0 Then a^2 = ab a^2 - b^2 = ab - b^2 (a + b)(a - b) = b(a - b) a + b = b b + b = b 2b = b 2 = 1 2 - 1 = 1 - 1 1 = 0 Granted, now we have 2 + 2 = 4 2 + 2 + 0 = 4 + 0 But since 1 = 0 2 + 2 + 0 = 4 + 1 2 + 2 = 5 QED I'm going to hell for this :D

OpenStudy (anonymous):

thanks.

terenzreignz (terenzreignz):

Hey, at least spot the error in the proof...

OpenStudy (anonymous):

where

terenzreignz (terenzreignz):

You know for a fact that 2 + 2 = 5 is false, yet I somehow "proved" it. Something must be wrong with the proof... don't give up so quickly, look more carefully :D

OpenStudy (anonymous):

ya i know, this is not true in the universe

terenzreignz (terenzreignz):

What if I tell you that it is? And I just proved it. Prove me wrong ;)

OpenStudy (mayankdevnani):

Start with: -20 = -20 Which is the same as: 16-36 = 25-45 Which can also be expressed as: (2+2) 2 (9 X (2+2) = 52) 9 X 5 Add 81/4 to both sides: (2+2) 2 (9 X (2+2) + 81/4 = 52) 9 X 5 + 81/4 Rearrange the terms: ({2+2}) 9/2) 2 = (5-9/2) 2 Ergo: 2+2 - 9/2 = 5 Hence: 2 + 2 = 5 where X=multiply(*)

OpenStudy (mayankdevnani):

i think it is correct @sunnymony

OpenStudy (anonymous):

no r u right.

OpenStudy (mayankdevnani):

what do you mean?

terenzreignz (terenzreignz):

@sunnymony At some point we had (a + b)(a - b) = b(a - b) a + b = b Which arose by dividing both sides by (a - b) Can you see the problem there? @mayankdevnani I don't see how you could arrive at 2 + 2 = 5 from 2 + 2 - 9/2 = 5

OpenStudy (mayankdevnani):

Start with: -20 = -20 OK Which is the same as: 16-36 = 25-45 OK Which can also be expressed as: (2+2) 2 (9 X (2+2) = 52) 9 X 5 Must have been written: (2+2)² - 9x(2+2) = 5² - 9x5 Add 81/4 to both sides: (2+2) 2 (9 X (2+2) + 81/4 = 52) 9 X 5 + 81/4 Must have been written: (2+2)² - 9x(2+2) + 81/4 = 5² - 9x5 + 81/4 Rearrange the terms: ({2+2}) 9/2) 2 = (5-9/2) 2 Must have been written: ((2+2) - 9/2)² = (5 - 9/2)² [step a] Ergo: 2+2 - 9/2 = 5 Must have been written: (2+2) - 9/2 = 5 - 9/2 [step b] Hence: 2 + 2 = 5

OpenStudy (mayankdevnani):

ok @sunnymony

OpenStudy (anonymous):

ok but one thing this is true in the universe

OpenStudy (mayankdevnani):

no it's just a math joke but don't take it simple....